OpenStudy (anonymous):

What's 2 x 6 - 3 + 9 ?

OpenStudy (anonymous):

First you need do PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

So 2 • 6 - 3 + 9 There is no Parentheses and Exponents, but there is Multiplication

OpenStudy (anonymous):

solve 2 • 6 and it was 12 12 - 3 + 9

OpenStudy (tyteen4a03):

2 * 6 = 12, 12 - 3 = 9, 9 + 9 = 18. Please use Wolfram|Alpha (or heck, even google) to do these simple calculations.

OpenStudy (anonymous):

then next is subtract 12 - 3 and it was 9 9 + 9 and it was 18

OpenStudy (anonymous):

Subtraction is first not addition because subtraction/or addition needs solve by left to right.